OpenStudy (anonymous):

Solve using row operations: x+2y=10; 3x+4y=8

OpenStudy (anonymous):

Subtract 2 times the 2nd row from the 1st row to get: x + 0 = -12 0 + y = 11

OpenStudy (michele_laino):

please multiply both sides of your first equation, namely x+2y=10, by -3, first